# Extending Quantum-Safe Communications to Real-World Networks: An Adaptive Security Framework

### v1

@@ IV-B Operational description of the security framework
Fig. 6 depicts the operational flow for the key derivation in level 4, for a configuration where the QuSeC configured a policy with dual KEM enforcement.
V VALIDATION
@@
This section presents the implementation used to validate and evaluate the performance of the proposed adaptive security framework. First, the testbed and implementation details are introduced. Next, the functional validation of the proposed security framework is presented, demonstrating its correct operation. Finally, performance evaluation results are discussed.
V-A Implementation and Testbed Setup
@@
The proposed solution was deployed on an experimental infrastructure consisting of several servers forming a virtualized node managed by an OpenStack orchestrator, providing more than 200 vCPUs, 500 GB of RAM, and 20 TB of storage. On this infrastructure, an experimental testbed that replicates the heterogeneous network topology depicted in Fig. 2 was deployed using Ubuntu 24.04 and Kubernetes orchestration. The testbed comprises six nodes: three Classical Nodes (CN A, B, C) and three QKD-enabled Nodes (QN D, E, F).
@@
  • Each vKMS instance supports the complete ETSI GS QKD 014 API to interface with end applications, extended with mechanisms for communication with the QuSeC to receive Security Levels and the related instructions. They support adaptive security mechanisms, including PQC integration and hybrid key derivation, under policy-driven Security Level execution. Post-quantum operations were realized with standardized algorithms, specifically CRYSTALS-Kyber (ML-KEM) [FIPS203] for KEM operations.
@@
  • The KMS instances also implement the ETSI GS QKD 014 API for interaction with the vKMSs for QKD key delivery, as well as the ETSI GS QKD 020 API for KMS interoperability in multi-hop QKD scenarios. In addition, they support the operations proposed in our previous work for communication with the QuSeC and the vKMS.
@@
  • The QuSeC implements the Security Level Assignment Function and associated configuration tasks, including topology aware path computation for multi-hop QKD scenarios, and automated policy generation for each Security Level.
@@
  • The QKD modules of the quantum plane have been implemented using SimulaQron, a distributed quantum network simulator designed for software development. However, as the framework relies on standardized ETSI APIs, it maintains full compatibility with real QKD hardware from any vendor. The simulator serves purely as a proof-of-concept platform, enabling comprehensive validation without requiring physical infrastructure.
V-B Functional validation
Comprehensive functional testing was carried out to validate the correct assignment of Security Levels across all communication scenarios. Table I summarizes the test cases executed, covering the four Security Levels. All test cases achieved 100% accuracy in Security Level Assignment, with the QuSeC correctly analyzing node capabilities and quantum path availability to determine the optimal configuration.
TABLE I: Test cases
Test Source App Target App Expected Level Assigned Level Result
T1 APP_E APP_F Level 1 Level 1 Pass
T2 APP_D APP_F Level 2 Level 2 Pass
T3 APP_C APP_D Level 3 Level 3 Pass
T4 APP_A APP_B Level 4 Level 4 Pass
In addition, for each Security Level, the complete E2E communication flow was validated, from initial application request to final key delivery.
  • Level 1 Validation: Direct QKD key retrieval between adjacent QKD nodes (E-F) completed successfully with information-theoretic security preservation. The vKMS instances correctly interfaced with the appropriate KMSs, KMS_E2 and KMS_F, via the ETSI GS QKD 014 API.
  • Level 2 Validation: Multi-hop QKD communication (D-F through trusted-relay node E) demonstrated proper relay path establishment and OTP-based key forwarding. The KMSs properly implemented the ETSI GS QKD 020 API for KMS interoperability for the QKD key relay.
  • Level 3 Validation: Hybrid CN-to-QN communication (C-D) validated the correct execution of the relay and hybridization mechanism: the QuSeC configured the appropriate roles and policies (Section IV-B2), and the protocol followed the six-step procedure defined in Section IV-B3, achieving successful PQC KEM establishment, QKD key relay, and final hybrid key derivation via the configured KDF.
  • Level 4 Validation: Pure PQC communication (A-B) demonstrated correct PQC-only operation: the QuSeC provided the policy configuration, and both vKMS instances executed the PQC KEM successfully, ensuring quantum-resistant protection in the absence of QKD infrastructure.
Overall, the implementation demonstrated full interoperability across heterogeneous node types and seamless integration between QKD and PQC. All key exchanges maintained compliance with ETSI specifications, while extending functionality to support adaptive security mechanisms.
V-C Performance Assessment
@@
To evaluate the performance of the proposed framework, latency measurements were collected for each Security Level. First, complete E2E key establishment across all Security Levels was analyzed, which accounts for all operations required for both the initiating and target applications. The delay associated with transmitting the key identifier from the initiating to the target application was excluded, as it is application-specific and out of the scope of this validation.
@@
In addition to the overall E2E latency, a detailed analysis was performed for each individual phase, Security Level Assignment, Configuration, and Secure Key Derivation, in order to quantify their individual contributions to total latency. Each test case defined in Table I was executed 100 times to obtain statistically significant measurements. The generated boxplots have been calculated with 98 percentile and show the median (horizontal line), the mean (x), the box (top limit 75 percentile and bottom limit 25 percentile), and the “whiskers” (that contain the 98% of results).
@@
Fig. 7 shows the total delay for E2E key establishment. Level 1 demonstrates the fastest establishment, averaging 73.35 ms through direct QKD key retrieval, followed by Level 4 with 92.3 ms via purely PQC-based operations. In contrast, the more complex Security Levels exhibited higher latencies, with 170.38 ms for Level 2 and 154.97 ms for Level 3. Despite the performance differences across levels, all key establishment operations remained below 200 ms, confirming the framework’s suitability.
@@
Moving towards the first subprocess, Security Level Assignment was highly consistent across all scenarios, with assignment times ranging from 8 to 9 ms for both initiating and target applications. This uniformity demonstrates that QuSeC performs the decision-making process efficiently, contributing negligible overhead to total E2E time.
The Security Level Configuration phase for the initiating application, shown in Fig. 8, varied more significantly between levels. Level 2 exhibited the highest latency at 58.35 ms, due to the complexity of relay path computation and rule installation across multiple intermediate KMS nodes. This reflects the coordination complexity demands for multi-hop QKD scenarios. The remaining Security Levels maintain low configuration times, with average values ranging from 6 to 11 ms. These results indicate that configuration complexity in the proposed framework is primarily driven by QKD relay infrastructure rather than PQC operations.
@@
The final phase, Secure Key Derivation, showed the largest variation across Security Levels, as depicted in Fig. 9 for initiating applications. Level 3 exhibited the highest processing time, averaging 98.09 ms, the complexity of the six-step hybrid mechanism that includes PQC KEM establishment, QKD key relay, OTP operations, and KDF processing. Level 2 followed with 53.30 ms, primarily due to multi-hop relay operations and OTP-based key forwarding. Conversely, Level 4 and Level 1 achieved significantly lower derivation times, 31.31 ms and 17.47 ms respectively, owing to their streamlined processes.
@@
Throughout all phases, target application procedures consistently demonstrate lower delays than those of the initiating application. This is a direct result of the QuSeC’s session state management, which prevents redundant processing for subsequent endpoints. As represented in Fig. 10, the target application’s contribution to total E2E time remains stable across all Security Levels, at approximately 30-35 ms, while the initiating application’s contribution varies according to the operations required by each subprocess.
@@
VI CONCLUSIONS
This work addresses the need for quantum-safe communications in real-world heterogeneous networks by introducing an adaptive framework that integrates QKD and PQC under centralized management. Unlike existing approaches limited to fully QKD-capable environments, this proposal extends quantum protection to all nodes, ensuring robust protection regardless of infrastructure constraints. By defining four dynamic security levels, the system enables tailored protection, balancing information-theoretic guarantees with practical deployability. While the default behavior prioritizes maximum security, the framework can be extended to consider specific service requirements and can even deny service requests when minimum security requirements cannot be satisfied.
The implementation and validation results demonstrate the framework’s ability to deliver secure keys across all communication scenarios. All tested security levels achieved reliable E2E operation, while maintaining full compliance with existing ETSI specifications. Performance evaluation further confirmed the feasibility of the proposed approach, highlighting the potential of the framework for deployment in transitional infrastructures where QKD adoption is partial or incremental.
In summary, this work establishes a practical path toward hybrid quantum-safe networks. By unifying QKD and PQC into a standard-aware, adaptive model, the framework provides the flexibility needed to adopt evolving technologies.

### v2

@@ IV-B Operational description of the security framework
Figure 6 depicts the operational flow for the key derivation in level 4, for a configuration where the QuSeC configured a policy with dual KEM enforcement.
V Security analysis
This section presents an overview of the security analysis of the QuLore framework under a Dolev-Yao network model [7]: the adversary can eavesdrop, delay, replay, inject, and arbitrarily modify any message transmitted over the classical network. Cryptographic primitives are analyzed in the standard computational model, considering a possibly quantum probabilistic-polynomial time (PPT) adversary, and capturing its ability to break the Secure Key Derivation procedure. The security analysis relies on the following security assumptions:
  1. 1.
    Authenticated secure channels: All control/management and key-delivery communications between applications, vKMSs, KMSs, and the QuSeC are transmitted over mutually authenticated and confidential channels.
  2. 2.
    ITS for QKD keys: Each QKD link generates information-theoretically secure random keys under standard QKD assumptions, including an authenticated classical channel for QKD post-processing.
  3. 3.
    Security of PQC and KDF: PQC KEMs are Indistinguishable under Adaptive Chosen-Ciphertext Attacks (IND-CCA2) against classical/quantum adversaries. Similarly, the KDF is modeled as PRF-like on its inputs, generating an indistinguishable from uniform output as long as at least one input remains secret.
  4. 4.
    No compromised nodes: The threat model excludes malicious insiders and compromised nodes. The attacker does not access internal states of entities, cannot extract keys from memory, and cannot perform physical extraction attacks. Trusted relay nodes in Levels 2–3 (Levels 1 and 4 do not comprise the use of trusted relays) are inside a declared trust boundary.
  5. 5.
    Controller key isolation: Architecturally, session keys and intermediate key material are never transmitted to QuSeC, which only sees metadata, IDs, and policies. Even if QuSeC were compromised beyond the primary model, it cannot directly learn session keys from protocol messages.
Under these assumptions, the security analysis focuses on session-key confidentiality, captured via an indistinguishability game (IND-SK): after interacting with honest protocol instances over the network, the adversary selects a completed test session and must distinguish the real session key from a uniformly random string of equal length. The protocol achieves confidentiality if the adversary’s advantage is negligible in the security parameter.
For each Security Level, full game-based proofs and reductions are performed, leading to the confidentiality guarantees summarized in Table I. For completeness, short proof sketches and theorems are included in this section.
In Level 1, the established key is directly a QKD-generated secret $K_{QKD}$. For an adversary to distinguish the real session key from random is by either eavesdropping on QKD channels (information-theoretically prevented by QKD), breaking the KMS protection (violating trust assumption), or recovering from network traffic (not possible due to authenticated and encrypted channels). Therefore, since the adversary’s information about the key is bounded information-theoretically, no computational advantage allows the adversary to distinguish the real key from random data, hence:
$\mathsf{Adv}^{\mathsf{IND\text{-}SK}}_{\text{Level 1}}(\lambda)\leq\text{negl}(\lambda)$
In Level 2, the established key is a QKD-derived secret $K_{0}$ that is relayed hop-by-hop using OTP encryption under independent per-link QKD keys. An adversary observing the classical network only sees relay ciphertexts $C_{i}=K_{0}\oplus K_{QKD,link_{i}}$, which are uniformly random due to OTP when the pad is secret and fresh, so they reveal no information about $K_{0}$. Therefore, the only practical way for the adversary to distinguish the real E2E session key from random is to (i) eavesdrop on QKD keys (information-theoretically prevented by QKD), (ii) compromise a trusted relay or access internal KMS/vKMS memory (violates the trust-boundary assumption), or (iii) recover key material from network traffic (not possible because key transfers inside nodes use mutually authenticated confidential channels). Hence, the adversary’s information about the session key remains bounded information-theoretically, and its IND-SK advantage is negligible under the stated assumptions, leading to the following theorem proof:
$\mathsf{Adv}^{\mathsf{IND\text{-}SK}}_{\text{Level 2}}(\lambda)\leq\text{negl}(\lambda)$
In Level 3, the session key is derived as $K_{session}=KDF(K_{QKD}||K_{PQC})$, where $K_{PQC}$ comes from a PQC KEM exchange and $K_{QKD}$ is obtained from the QKD link and relayed to the CN using OTP under an auxiliary PQC-derived secret $K_{PQC,aux}$. From the adversary’s perspective, the only way to distinguish the real session key from random is to either (i) break the IND-CCA2 security of the KEM to recover $K_{PQC}$ (and potentially $K_{PQC,aux}$), (ii) distinguish the KDF output from random, breaking the PRF-like property of the KDF, (iii) violate the QKD secrecy assumptions or compromise internal KMS/vKMS state (out of scope), or (iv) recover secrets from protected network traffic (not possible under the secure-channel assumption). Therefore, any non-negligible adversary can be reduced to breaking IND-CCA2 (KEM) or PRF (KDF), so the IND-SK advantage is negligible under those security assumptions, as stated in the following theorem:
$\mathsf{Adv}^{\mathsf{IND\text{-}SK}}_{\text{Level 3}}(\lambda)\leq\mathsf{Adv}^{\text{IND-CCA2}}_{\text{KEM}}(\lambda)+\mathsf{Adv}^{\text{PRF}}_{\text{KDF}}(\lambda)+\text{negl}(\lambda)$
In Level 4, both endpoints are classical and establish the session key purely using PQC KEM, and optionally a KDF if combining multiple KEM outputs. For the single-KEM variant, an adversary can only distinguish the real session key from random by breaking the KEM’s IND-CCA2 security, or violating the protected-channel/trusted-node assumptions. Otherwise the KEM shared secret is computationally indistinguishable from uniform, and so is the derived session key. For the dual-KEM configuration, the adversary’s task is even harder: to distinguish the final key, it must break at least one of the assumed components: IND-CCA2 of KEM1, IND-CCA2 of KEM2, or PRF security of the KDF. Otherwise, the combined output remains indistinguishable from random. Therefore, under IND-CCA2 security of the employed KEM(s) and PRF-like KDF when used, the adversary’s distinguishing advantage in IND-SK is negligible. The following theorems show the advantage of an adversary for single KEM and dual KEM cases, respectively:
$\mathsf{Adv}^{\mathsf{IND\text{-}SK}}_{\text{Level 4 (single)}}(\lambda)\leq\mathsf{Adv}^{\text{IND-CCA2}}_{KEM}(\lambda)+\text{negl}(\lambda)$
$\mathsf{Adv}^{\mathsf{IND\text{-}SK}}_{\text{Level 4 (dual)}}(\lambda)\leq\mathsf{Adv}^{\text{IND-CCA2}}_{\text{KEM}_{1}}(\lambda)+\mathsf{Adv}^{\text{IND-CCA2}}_{\text{KEM}_{2}}(\lambda)$
$\mathsf{Adv}^{\text{PRF}}_{\text{KDF}}(\lambda)+\text{negl}(\lambda)$
TABLE I: Confidentiality guarantees per Security Level under the threat model.
Level Construction Confidentiality guarantee Main assumptions
L1 Direct QKD key delivery Information-theoretic secrecy QKD assumptions; protected key delivery channels
L2 Multi-hop relay using OTP on per-link QKD keys Information-theoretic secrecy within trust boundary Trusted relays (no compromise), fresh OTP keys per hop
L3 Hybrid: $K=KDF(K_{QKD}\,\|\,K_{PQC})$ with QKD relay protected by PQC-derived auxiliary secret Computational IND-SK confidentiality IND-CCA2(KEM), PRF(KDF), trusted relay
L4 PQC-only (single KEM or dual-KEM + KDF) Computational IND-SK confidentiality IND-CCA2(KEM), PRF(KDF)
VI Performance evaluation
@@
This section presents the implementation used to validate and evaluate the performance of the QuLore framework. First, the testbed and implementation details are introduced. Next, the functional validation of the proposed security framework is presented, demonstrating its correct operation. Finally, performance evaluation results are discussed.
VI-A Implementation and Testbed Setup
@@
The proposed solution was deployed on a Supermicro SYS-E200-8D server, featuring an Intel(R) Xeon(R) D-1528 processor running at 1.90 GHz on an x86_64 architecture with 6 cores, backed by 64 GB of RAM and approximately 500 GB of local storage. On this infrastructure, an experimental testbed that replicates the heterogeneous network topology depicted in Figure 2 was deployed using Ubuntu 20.04 and Kubernetes orchestration. The testbed comprises six nodes: three Classical Nodes (CN A, B, C) and three QKD-enabled Nodes (QN D, E, F).
@@
  • Each vKMS instance supports the complete ETSI GS QKD 014 API to interface with end applications, extended with mechanisms for communication with the QuSeC to receive Security Levels and the related instructions. They support adaptive security mechanisms, including PQC integration and hybrid key derivation, under policy-driven Security Level execution. Post-quantum operations were realized with standardized algorithms, specifically ML-KEM [1] for KEM operations.
@@
  • The KMS instances also implement the ETSI GS QKD 014 API for interaction with the vKMSs for QKD key delivery, as well as the ETSI GS QKD 020 API draft for KMS interoperability in multi-hop QKD scenarios. In addition, they support the operations proposed in our previous work for communication with the QuSeC and the vKMS.
@@
  • The QuSeC implements the Security Level Assignment Function and associated configuration tasks, including topology-awaree path computation for multi-hop QKD scenarios, and automated policy generation for each Security Level. The path computation assumes equal link costs, so the routing criterion corresponds to the minimum hop count, although these weights could be dynamically adjusted to reflect additional metrics such as available key rate.
@@
  • The QKD modules of the quantum plane have been implemented using SimulaQron, a distributed quantum network simulator designed for software development. However, as the framework relies on standardized ETSI APIs, it maintains full compatibility with real QKD hardware from any vendor. The simulator serves purely as a proof-of-concept platform, enabling comprehensive validation without requiring physical infrastructure. Performance validation in this work targets the proposed framework, which is independent of the underlying quantum-channel key rate. Therefore, SimulaQron is used solely as a key source to populate the KMSs, and can be transparently replaced by real QKD hardware without affecting the reported performance results.
Finally, to ensure consistency and reproducibility, all cryptographic operations are configured with standardized parameters. Table II summarizes the main implementation settings, including the 256-bit key size for all cryptographic mechanisms and the use of different algorithms.
TABLE II: Implementation details.
Component Details
Key length 256 bits
[…]
PQC KEM algorithm ML-KEM-768
KDF algorithm HKDF-SHA256
Software implementation with SimulaQron simulator.
VI-B Definition of Key Performance Indicators
Based on the QuLore framework implementation and the deployed testbed, to assess the feasibility and performance of QuLore, different KPIs were defined:
  1. 1.
    Latency of the E2E key establishment procedure for each Security Level.
  2. 2.
    Latency of each main subprocess of the framework for each Security Level:
    1. (a)
      Latency of Security Level Assignment.
    2. (b)
      Latency of Security Level Configuration.
    3. (c)
      Latency of Secure Key Derivation.
  3. 3.
    CPU usage during the E2E key establishment for each Security Level.
  4. 4.
    Overhead of the communication associated with each Security Level.
The following subsection presents and discusses the experimental results, providing insights into the latency behaviour, CPU consumption, and communication overhead introduced by each process.
VI-C Results and Discussion
@@
To evaluate the performance of QuLore, latency measurements were collected for each Security Level. First, complete E2E key establishment across all Security Levels was analyzed, which accounts for all operations required for both the initiating and target applications. The delay associated with transmitting the key identifier from the initiating to the target application was excluded, as it is application-specific and out of the scope of this validation.
@@
In addition to the overall E2E latency, a detailed analysis was performed for each individual phase, Security Level Assignment, Configuration, and Secure Key Derivation, in order to quantify their individual contributions to total latency. Each test case defined in Table III (same as those defined in Figure 4) was executed 100 times to obtain statistically significant measurements. The generated boxplots have been calculated with 98 percentile and show the median (horizontal line), the mean (x), the box (top limit 75 percentile and bottom limit 25 percentile), and the “whiskers” (that contain the 98% of results).
Beyond latency, CPU consumption was measured using the resource package in Python, which captures the cumulative user and system time consumed by each process during each execution. Communication overhead was assessed by analyzing the communication traces generated for each test case, enabling quantification of the protocol-induced overhead independent of application-level data exchange.
TABLE III: Test cases
Test Source App Target App Security Level
T1 APP_E APP_F Level 1
T2 APP_D APP_F Level 2
T3 APP_C APP_D Level 3
T4 APP_A APP_B Level 4
@@
Figure 7 shows the total delay for E2E key establishment. Level 1 demonstrates the fastest establishment, averaging 59.81 ms through direct QKD key retrieval, followed by Level 4 with 85.06 ms via purely PQC-based operations. In contrast, the more complex Security Levels exhibited higher latencies, with 138.92 ms for Level 2 and 143.27 ms for Level 3. Additionally, this latency is highly dependent on the number of intermediate trusted nodes that are involved in the key establishmentment. Despite the performance differences across levels, all key establishment operations remained below 150 ms, confirming the framework’s suitability.
@@
Moving towards the first subprocess, Security Level Assignment was highly consistent across all scenarios, with assignment times around 6 ms for both initiating and target applications. This uniformity demonstrates that QuSeC performs the decision-making process efficiently, contributing negligible overhead to total E2E time.
The Security Level Configuration phase for the initiating application, shown in Figure 8, varied more significantly between levels. Level 2 exhibited the highest latency at 39.19 ms, due to the complexity of relay path computation and rule installation across multiple intermediate KMS nodes. This reflects the coordination complexity demands for multi-hop QKD scenarios. The remaining Security Levels maintain low configuration times, with average values ranging from 5 to 7 ms. These results indicate that configuration complexity in the proposed framework is primarily driven by QKD relay infrastructure rather than PQC operations.
@@
The final phase, Secure Key Derivation, showed the largest variation across Security Levels, as depicted in Figure 9 for initiating applications. Level 3 exhibited the highest processing time, averaging 88.63 ms, the complexity of the six-step hybrid mechanism that includes PQC KEM establishment, QKD key relay, OTP operations, and KDF processing. Level 2 followed with 54.10 ms, primarily due to multi-hop relay operations and OTP-based key forwarding. Conversely, Level 4 and Level 1 achieved significantly lower derivation times, 31.47 ms and 11.38 ms respectively, owing to their streamlined processes.
@@
Throughout all phases, target application procedures consistently demonstrate lower delays than those of the initiating application. This is a direct result of the QuSeC’s session state management, which prevents redundant processing for subsequent endpoints. As represented in Figure 10, the target application’s contribution to total E2E time remains stable across all Security Levels, at approximately 30-35 ms, while the initiating application’s contribution varies according to the operations required by each subprocess.
@@
Regarding CPU consumption during E2E key establishment, results show that CPU cost increases with protocol complexity across Security Levels and remains within the sub-120 ms range in the evaluated setup: 53.43 ms (L1), 104.35 ms (L2), 77.10 ms (L3), and 54.66 ms (L4) of total CPU time per establishment, as depicted in Figure 11.
Figure 11: Total CPU usage of the E2E key establishment across levels.
Beyond absolute CPU time, the ratio between CPU time and E2E latency indicates whether performance is dominated by computation or by distributed coordination. Level 1 is largely CPU-bound (53.43 ms CPU vs 59.81 ms latency, 90%), reflecting a short workflow with limited messaging. Level 2 remains mostly CPU-bound but with a larger coordination component (104.35 ms CPU vs 138.92 ms latency, 75%), consistent with additional processing required for trusted-relay handling. In contrast, Level 3 is primarily coordination/communication-bound (77.10 ms CPU vs 143.27 ms latency,  54%), since its hybrid procedure introduces multiple message exchanges and waiting periods even though local computation remains moderate. Level 4 shows intermediate behavior (54.66 ms CPU vs 85.06 ms latency,  65%), where the remaining gap is attributable to communication and protocol orchestration.
Finally, communication overhead was quantified as the total number of bytes exchanged during E2E establishment, with results summarized in Figure 12. Overhead remains small across all Security Levels: 3.46 kB (L1), 7.25 kB (L2), 9.94 kB (L3), and 6.02 kB (L4). Even the most expensive configuration (L3) stays below 10 kB per key establishment, indicating that the protocol is lightweight in terms of network load. It is important to note that this overhead remains consistent across all Security Levels, except for Level 2, where additional hops require extra path installation rules and key relay messages.
Figure 12: Communication overhead of the E2E key establishment across Security Levels.
VII Conclusions
This work addresses the need for quantum-safe communications in real-world heterogeneous networks by introducing QuLore: an adaptive framework that integrates QKD and PQC under centralized management. Unlike existing approaches limited to fully QKD-capable environments, this proposal extends quantum protection to all nodes, ensuring robust protection regardless of infrastructure constraints. By defining four dynamic security levels, the system enables tailored protection, balancing information-theoretic guarantees with practical deployability. While the default behavior prioritizes maximum security, the framework can be extended to consider specific service requirements and can even deny service requests when minimum security requirements cannot be satisfied.
The implementation and validation results demonstrate the QuLore framework’s ability to deliver secure keys across all communication scenarios in a secure manner. All tested security levels achieved reliable E2E operation, while maintaining full compliance with existing ETSI specifications. Performance evaluation further confirmed the feasibility of the proposed approach, highlighting the potential of the framework for deployment in transitional scenarios where QKD adoption is partial or incremental. Additionally, while the centralized design introduces a potential single point of failure, this risk can be mitigated by deploying QuSeC in a replicated configuration with state synchronization across redundant instances via east–west interfaces. This approach also supports scalability in large deployments by distributing control-plane load without affecting key confidentiality.
In summary, this work establishes a practical path toward hybrid quantum-safe networks. By unifying QKD and PQC into a standard-aware, adaptive model, QuLore provides the flexibility needed to adopt evolving technologies.